# ScrewSplat: An End-to-End Method for Articulated Object Recognition

### v1

@@ Abstract
Articulated object recognition – the task of identifying both the geometry and kinematic joints of objects with movable parts – is essential for enabling robots to interact with everyday objects such as doors and laptops. However, existing approaches often rely on strong assumptions, such as a known number of articulated parts; require additional inputs, such as depth images; or involve complex intermediate steps that can introduce potential errors – limiting their practicality in real-world settings. In this paper, we introduce ScrewSplat, a simple end-to-end method that operates solely on RGB observations. Our approach begins by randomly initializing screw axes, which are then iteratively optimized to recover the object’s underlying kinematic structure. By integrating with Gaussian Splatting, we simultaneously reconstruct the 3D geometry and segment the object into rigid, movable parts. We demonstrate that our method achieves state-of-the-art recognition accuracy across a diverse set of articulated objects, and further enables zero-shot, text-guided manipulation using the recovered kinematic model.
@@ 1 Introduction
Spurred in part by the success of neural rendering-based 3D reconstruction methods that require no supervised training [12, 13, 14, 15], recent works have adapted these frameworks for articulated object recognition [16, 17, 18, 19, 20], achieving promising results using only raw RGB observations. However, a key drawback of these methods lies in their reliance on strong assumptions, such as a known number of articulated components or predefined joint types. Moreover, they often involve multi-stage pipelines with intermediate procedures like point correspondence matching or part clustering, which not only increase overall complexity but also introduce potential points of failure – errors in these stages can significantly impair final performance. Furthermore, some of these approaches rely on auxiliary depth inputs, which are often noisier than RGB images – particularly for transparent or reflective surfaces – thereby limiting their robustness in real-world scenarios.
We propose ScrewSplat, a simple end-to-end method for articulated object recognition that avoids intermediate steps, auxiliary data, and prior knowledge of joint types or counts. We formulate the task as joint optimization over part-aware geometry, joint axes and types, and joint angles, such that the rendered views match the observations. This problem is particularly challenging due to its hybrid structure – continuous variables (geometry and joint angles) are coupled with discrete, combinatorial ones (part segmentation labels, joint types, and joint counts).
@@ 2 Related Works
Supervised Learning from Articulated Object Datasets. Earlier works on articulated object recognition primarily focus on supervised methods, which train deep neural networks to predict articulation parameters (e.g., joint axes and joint angles) directly from raw visual input [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. Some of these methods also aim to simultaneously reconstruct the part-aware geometry of objects [5, 8, 10], which aligns with our objectives. While they demonstrate generalizability to unseen objects within similar categories, they often struggle to generalize to objects that differ significantly from the training categories.
Per-object Optimization Methods. To circumvent the need for explicit supervision, several works have attempted to recognize articulated objects by directly fitting 3D representations – such as neural rendering methods [12, 13, 14, 15] – and kinematic structures to RGB observations [16, 17, 18, 19, 20]. Due to their reliance on object-specific optimization, these approaches are often referred to as per-object optimization methods. While these methods demonstrate strong performance in a category-agnostic setting, they suffer from a critical limitation: they rely on strong assumptions, such as a known number of articulated joints or even predefined articulation types. This limitation stems from their typical decomposition of the problem into discovering the static and movable parts of an object, followed by analyzing the motion of the movable parts [17, 18, 19, 20]. Accurate part discovery generally requires prior knowledge of the number of parts to enable meaningful segmentation, which limits the applicability of these methods in real-world scenarios.
@@ 3 Preliminaries
In this section, we introduce two key components that form the foundation of our approach. First, we provide a brief overview of screw theory – a fundamental concept in robotics and rigid body kinematics – that enables efficient modeling of joint motion and spatial transformations. Next, we review 3D Gaussian splatting, a recently introduced technique for representing and rendering 3D scenes using anisotropic Gaussian primitives.
@@ 3.1 Screw Theory
[…]
@@
where $T^{\prime}\in\mathrm{SE}(3)$ denotes the transformed rigid body coordinate, and $[\mathcal{S}]$ is the $\times 4$ 4 matrix representation of the screw axis $\mathcal{S}$, defined as
$[\mathcal{S}]=\begin{bmatrix}[\omega]&v\\ 0&0\end{bmatrix}\in\mathbb{R}^{4\times 4},\>\>\>[\omega]=\begin{bmatrix}0&-\omega_{3}&\omega_{2}\\ \omega_{3}&0&-\omega_{1}\\ -\omega_{2}&\omega_{1}&0\end{bmatrix}\in\mathbb{R}^{3\times 3},$ (3)
where $\omega=(\omega_{1},\omega_{2},\omega_{3})$. Figure 2 illustrates the revolute and prismatic screw axes and their corresponding screw motions. The matrix exponential formulation admits closed-form expressions for both revolute and prismatic screw axes; further details are provided in Appendix B.
@@ 4.1 ScrewSplat
The optimization procedure generally follows that of the original Gaussian Splatting, with a few additional modifications. The most important modification is that all screw confidences $\gamma_{j}$ and part probabilities $m_{i}$ are periodically reset to 0.9 and uniform distribution, respectively. This periodic reinitialization helps ScrewSplat effectively discover meaningful screw primitives. After optimization, only the screw primitives that satisfy certain criteria remain, ensuring that only the most relevant ones are retained. Further optimization details can be found in Appendix B.
@@ 4.2 Controlling Joint Angles Using ScrewSplat as a Renderer
$\mathcal{L}_{\text{CLIP-dir}}=1-\frac{\triangle I(\theta)\cdot\triangle T}{||\triangle I(\theta)||\>||\triangle T||},$ (8)
@@
After obtaining the target joint angle vector $\theta$, we plan a simple trajectory for the robot end-effector tip and generate a kinematically feasible robot trajectory accordingly, as shown on the right of Figure 4. Further details of the optimization and planning procedure are provided in Appendix B.
@@ 5 Experiments
Table 1: Recognition performance for single-joint objects, averaged across all instances. Object-wise recognition performance is provided in Appendix D.
@@
Baseline Methods. We compare ScrewSplat with the following baselines: PARIS [16], PARIS* [16], DTA [17], and ScrewSpawn. PARIS and PARIS* (PARIS augmented with depth data) recover a single joint axis assuming a known joint type, while DTA discovers a predefined number of joint axes. Since only DTA supports multiple joint axes, we compare against DTA for multi-joint objects. These methods recover kinematic structures from multi-view images under two different object configurations; PARIS uses RGB input, while PARIS* and DTA use RGB-D. We also introduce ScrewSpawn, an ablation model that follows the ScrewSplat framework but spawns only a single screw (with a known joint type). This model is used to validate the necessity of “splatting” multiple screw primitives in ScrewSplat. Detailed implementations can be found in Appendix C.
Dataset. We select ten single-joint objects and three multi-joint objects from distinct categories in the PartNet-Mobility dataset [25]. Using Blender [26], a photorealistic renderer, we obtain multi-view RGB images and depth images (used for training PARIS* and DTA) under varying object configurations. We place 48 camera poses uniformly over a hemisphere centered on each object. For ScrewSplat and ScrewSpawn, we use images captured under five configurations, while for PARIS, PARIS*, and DTA, we use images from two configurations selected from the same five. For single-joint objects, the five configurations correspond to evenly spaced joint angles along the motion range. For multi-joint objects, five randomly sampled joint angle vectors are used. Further details for the evaluation dataset are provided in Appendix C.
@@ 5.1 Articulated Object Recognition Performance
[…]
@@
Single-joint Objects. Figure 5 shows the recognition results, including reconstructed meshes and screw axes, on three representative single-joint objects. PARIS and PARIS* performed well on the folding chair but failed to recognize the other objects. DTA generally outperforms PARIS, but still fails in certain cases such as the stapler. ScrewSplat consistently demonstrates the best performance across all cases, both in terms of geometry reconstruction and kinematic structure estimation. In contrast, ScrewSpawn fails to accurately recover both geometry and kinematic structure in most cases, highlighting the effectiveness of ScrewSplat’s multi-screw formulation. Table 1 shows the quantitative recognition results. We demonstrate that ScrewSplat achieves the highest overall performance across geometry, motion, and visual appearance. PARIS and PARIS* struggle to recognize almost all of the other objects. ScrewSpawn is able to reconstruct the static parts to some extent but fails to accurately estimate the movable parts and screw axes. DTA performs slightly worse than our method but achieves comparable results overall. Although our method uses data from a wider range of joint configurations, it is particularly notable that it achieves the best performance while relying solely on RGB inputs and without any prior knowledge of the joints. Detailed object-wise recognition results are provided in Appendix D.
@@
Table 2: Recognition performance on multi-joint objects, averaged over all instances, geometries, and screw axes. Detailed object-wise recognition results are provided in Appendix D.
@@
METHOD CD-s CD-mm CD-w Angm. Posm. PSNR SSIM
@@ 5.2 Articulated Object Manipulation Results
Real-world Articulated Object Manipulation. For real-world scenarios where a robot manipulator physically adjust the joint angles of articulated objects, the overall manipulation pipeline consists of three main stages, as illustrated in Figure 8: (i) a recognition step, where ScrewSplat is optimized using multi-view RGB observations collected under several joint configurations manually manipulated by a human; (ii) a current state estimation step, which estimates the object’s current joint angle by optimizing an appropriate loss function; and (iii) a text-guided manipulation step, which determines the target joint angle based on a given text prompt and executes the corresponding robot manipulation. As shown on the left of Figure 8, ScrewSplat accurately reconstructs the shape and kinematic structure of real-world objects – including even a translucent drawer. A well-trained ScrewSplat further enables precise estimation of the current joint angle and facilitates successful text-guided object manipulation, as shown in the middle and right of Figure 8, respectively. Further details for real-world text-guided manipulation are provided in Appendix C and D, respectively.
@@ Acknowledgments
S. Kim, J. Ha, Y. H. Kim, and F. C. Park were supported in part by IITP-MSIT under Grant RS-2021-II212068 (SNU AI Innovation Hub), in part by IITP-MSIT under Grant 2022-220480 and Grant RS-2022-II220480 (Training and Inference Methods for Goal Oriented AI Agents), in part by IITP-MSIT under Grant RS-2024-00436680 (Collaborative Research Projects with Microsoft Research) through Global Research Support Program in the Digital Field program, in part by KIAT under Grant P0020536 (HRD Program for Industrial Innovation), in part by SRRC NRF under Grant RS-2023-00208052, in part by SNU-AIIS, in part by SNU-IPAI, in part by SNU-IAMD, in part by SNU BK21+ Program in Mechanical Engineering, in part by SNU Institute for Engineering Research, in part by Hyundai Motor Company and Kia, and in part by Microsoft Research Asia.
@@ A.1 Articulated Object Recognition Using Supervised Learning
There has been considerable interest in recognizing articulated objects within a supervised learning framework, where deep neural networks are trained to predict the underlying kinematic structure directly from raw visual input. Early studies often addressed this by employing category-driven approaches, using category information to assist in identifying the kinematic structure, and testing on unseen objects belonging to known categories [1, 4, 7, 32]. More recently, to handle a broader range of objects beyond those in known categories, category-agnostic recognition approaches have been proposed. These methods primarily aim to reconstruct part-level geometry and kinematic structures, which closely align with our objective. For example, CARTO [10] predicts the geometry as a defomable signed distance function (SDF) representations – similar to the category-specific A-SDF [4] but removes the reliance on object categories. Ditto [8] predicts articulable neural occupancy fields to generate digital twins of articulated objects from point cloud observations under two different articulation states. Articulate-Anything [33], a recent approach, integrates a vision-language foundation model to recognize part-aware geometries and kinematic structures, and incorporates interactable digital twins into simulators for sim-to-real robot learning. They demonstrate generalizability to unseen objects within similar categories, and even to unseen categories through foundation models. However, they inherently struggle to generalize to objects that differ significantly from the training categories.
@@ A.2 Articulated Object Recognition Using Per-object Optimization Methods
Several works have attempted to recognize articulated objects by directly fitting 3D representations and kinematic structures to observations without any supervision. Since these methods typically perform optimization for each object individually, they are often referred to as per-object optimization methods. A representative early work in per-object optimization methods is PARIS [16], which presents a method based on neural radiance fields (NeRF). Specifically, PARIS defines separate radiance fields for movable and static parts, performing joint rendering and optimization to achieve part-level reconstruction and articulation discovery. However, this approach is only applicable to single-joint objects, i.e., articulated objects with only one movable part, with pre-determined joint types.

### v2

@@ Abstract
Articulated object recognition – the task of identifying both the geometry and kinematic joints of objects with movable parts – is essential for enabling robots to interact with everyday objects such as doors and laptops. However, existing approaches often rely on strong assumptions, such as a known number of articulated parts; require additional inputs, such as depth images; or involve complex intermediate steps that can introduce potential errors – limiting their practicality in real-world settings. In this paper, we introduce ScrewSplat, a simple end-to-end method that operates solely on RGB observations. Our approach begins by randomly initializing screw axes, which are then iteratively optimized to recover the object’s underlying kinematic structure. By integrating with Gaussian Splatting, we simultaneously reconstruct the 3D geometry and segment the object into rigid, movable parts. We demonstrate that our method achieves state-of-the-art recognition accuracy across a diverse set of articulated objects, and further enables zero-shot, text-guided manipulation using the recovered kinematic model. See the project website at: https://screwsplat.github.io.
@@ 1 Introduction
Spurred in part by the success of neural rendering-based 3D reconstruction methods that require no supervised training [12, 13, 14, 15], recent works have adapted these frameworks for articulated object recognition [16, 17, 18, 19, 20], achieving promising results using raw RGB observations. However, a key drawback of these methods lies in their reliance on strong assumptions, such as a known number of articulated components or predefined joint types. Moreover, they often involve multi-stage pipelines with intermediate procedures like point correspondence matching or part clustering, which not only increase overall complexity but also introduce potential points of failure – errors in these stages can significantly impair final performance. Furthermore, some of these approaches rely on auxiliary depth inputs, which are often noisier than RGB images – particularly for transparent or reflective surfaces – thereby limiting their robustness in real-world scenarios.
We propose ScrewSplat, a simple end-to-end method for articulated object recognition that avoids intermediate steps, auxiliary data, and any prior knowledge of joint types or counts. We formulate the task as a joint optimization over part-aware geometry, joint axes and types, and joint angles, such that the rendered views match the observations. This problem is particularly challenging due to its hybrid structure – continuous variables (geometry and joint angles) are coupled with discrete, combinatorial ones (part segmentation labels, joint types, and joint counts).
@@ 2 Related Works
This section focuses on articulated object recognition methods that do not rely on explicit supervision [16, 17, 18, 19, 20]. A detailed review of supervised methods is provided in Appendix A.1.
A representative early work is PARIS [16], which achieves part-level reconstruction and articulation discovery with NeRF, but is restricted to single-joint objects with known joint types. DTA [17], a more recent approach, extend to multi-joint objects by reconstructing meshes from RGB-D data and inferring kinematic structures via correspondence matching; however, they still depend on depth input and prior knowledge of the number of movable parts. Gaussian splatting-based methods, including ArtGS [19] and ArticulatedGS [20], provide an alternative. While ArtGS maintains assumptions similar to DTA, ArticulatedGS relaxes them, but remains limited to recovering only one articulation per optimization step. A more detailed review of these approaches is provided in Appendix A.2.
Collectively, while these methods demonstrate strong performance in a category-agnostic setting, they suffer from a critical limitation – reliance on strong assumptions, such as prior knowledge of the number of articulated joints or even predefined articulation types. This limitation arises from their common strategy of decomposing the problem into the discovery of static and movable object parts, followed by analyzing the motion of the movable parts [17, 18, 19, 20]. Accurate part discovery generally requires prior knowledge of the number of parts to achieve meaningful segmentation, which restricts the applicability of these methods in real-world scenarios.
@@ 3 Preliminaries
In this section, we introduce two core components of our approach. First, we briefly review screw theory – a fundamental concept in robotics and rigid-body kinematics that enables efficient modeling of joint motion and spatial transformations. Next, we outline 3D Gaussian splatting, a recent technique for representing and rendering 3D scenes with anisotropic Gaussian primitives.
@@ 3.1 Screw Theory
where $\omega\in\mathbb{R}^{3}$ and $v\in\mathbb{R}^{3}$, and they satisfy either (i) $\omega||=1$ or (ii) $\omega=0$ and v=1||v||=1.
@@
where $T^{\prime}\in\mathrm{SE}(3)$ denotes the transformed rigid body coordinate, and $\mathcal{S}]$ is the $4\times 4$ matrix representation of the screw axis $\mathcal{S}$, defined as
$\mathcal{S}]=\begin{bmatrix}[\omega]&v\\ 0&0\end{bmatrix}\in\mathbb{R}^{4\times 4},\>\>\>[\omega]=\begin{bmatrix}0&-\omega_{3}&\omega_{2}\\ \omega_{3}&0&-\omega_{1}\\ -\omega_{2}&\omega_{1}&0\end{bmatrix}\in\mathbb{R}^{3\times 3},$ (3)
where $\omega=(\omega_{1},\omega_{2},\omega_{3})$. Figure 2 illustrates the revolute and prismatic screw axes and their corresponding screw motions. The matrix exponential formulation admits closed-form expressions for both revolute and prismatic screw axes; further details are provided in Appendix B.1.
@@ 4.1 ScrewSplat
The optimization procedure generally follows that of the original Gaussian Splatting, with a few additional modifications. The most important modification is that all screw confidences $\gamma_{j}$ and part probabilities $m_{i}$ are periodically reset to 0.9 and uniform distribution, respectively. This periodic reinitialization helps ScrewSplat effectively discover meaningful screw primitives. After optimization, only the screw primitives that satisfy certain criteria remain, ensuring that only the most relevant ones are retained. Further optimization details can be found in Appendix B.2.
@@ 4.2 Controlling Joint Angles Using ScrewSplat as a Renderer
[…]
@@
After obtaining the target joint angle vector $\theta$, we plan a simple trajectory for the robot end-effector tip and generate a kinematically feasible robot trajectory accordingly, as shown on the right of Figure 4. Further details of the optimization and planning procedure are provided in Appendix B.3.
@@ 5 Experiments
Table 1: Recognition performance for single-joint objects, averaged across all instances. Object-wise recognition performance is provided in Appendix D.1.
@@
Baseline Methods. We compare ScrewSplat with the following baselines: PARIS [16], PARIS*, DTA [17], and ScrewSpawn. PARIS and PARIS* (PARIS augmented with depth data) recover a single joint axis assuming a known joint type, while DTA discovers a predefined number of joint axes. Since only DTA supports multiple joint axes, we compare against DTA for multi-joint objects. These methods recover kinematic structures from multi-view images under two different object configurations; PARIS uses RGB input, while PARIS* and DTA use RGB-D. We also introduce ScrewSpawn, an ablation model that follows the ScrewSplat framework but spawns only a single screw (with a known joint type). This model is used to validate the necessity of “splatting” multiple screw primitives in ScrewSplat. Detailed implementations can be found in Appendix C.1.
Dataset. We select ten single-joint objects and three multi-joint objects from distinct categories in the PartNet-Mobility dataset [25]. Using Blender [26], a photorealistic renderer, we obtain multi-view RGB images and depth images (used for training PARIS* and DTA) under varying object configurations. We place 48 camera poses uniformly over a hemisphere centered on each object. For ScrewSplat and ScrewSpawn, we use images captured under five configurations, while for PARIS, PARIS*, and DTA, we use images from two configurations selected from the same five. For single-joint objects, the five configurations correspond to evenly spaced joint angles along the motion range. For multi-joint objects, five randomly sampled joint angle vectors are used. Further details for the evaluation dataset are provided in Appendix C.1.
@@ 5.1 Articulated Object Recognition Performance
We compare the performance of ScrewSplat with the baselines. To evaluate the recognition quality of articulated objects, we adopt three types of metrics: geometry, motion, and appearance. The geometry metric includes the bi-directional Chamfer-$l2$ distance between point clouds sampled from the reconstructed and ground-truth meshes. We report this metric separately for the static part (CD-s, mm), the movable parts (CD-m, mm), and the entire object (CD-w, mm). The motion metric includes the angular error (Ang., ) between the estimated and ground-truth screw axes, and the axis position error (Pos., 0.1m) for revolute joints, calculated as the minimum distance between corresponding screw axes. The appearance metric includes Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM), computed on rendered images under unseen joint angles. Details of the evaluation metrics and their computation are available in Appendix C.1.
@@
Single-joint Objects. Figure 5 shows the recognition results, including reconstructed meshes and screw axes, on three representative single-joint objects. PARIS and PARIS* performed well on the folding chair but failed to recognize the other objects. DTA generally outperforms PARIS, but still fails in certain cases such as the stapler. ScrewSplat consistently demonstrates the best performance across all cases, both in terms of geometry reconstruction and kinematic structure estimation. In contrast, ScrewSpawn fails to accurately recover both geometry and kinematic structure in most cases, highlighting the effectiveness of ScrewSplat’s multi-screw formulation. Table 1 shows the quantitative recognition results. We demonstrate that ScrewSplat achieves the highest overall performance across geometry, motion, and visual appearance. PARIS and PARIS* struggle to recognize almost all of the other objects. ScrewSpawn is able to reconstruct the static parts to some extent but fails to accurately estimate the movable parts and screw axes. DTA performs slightly worse than our method but achieves comparable results overall. Although our method uses data from a wider range of joint configurations, it is particularly notable that it achieves the best performance while relying solely on RGB inputs and without any prior knowledge of the joints. Detailed object-wise results are provided in Appendix D.1.
@@
Table 2: Recognition performance on multi-joint objects, averaged over all instances, geometries, and screw axes. Detailed object-wise recognition results are provided in Appendix D.1.
@@
METHOD CD-s CD-$m$ CD-w Angm. Posm. PSNR SSIM
@@ 5.2 Articulated Object Manipulation Results
Real-world Articulated Object Manipulation. For real-world scenarios where a robot manipulator physically adjust the joint angles of articulated objects, the overall manipulation pipeline consists of three main stages, as illustrated in Figure 8: (i) a recognition step, where ScrewSplat is optimized using multi-view RGB observations collected under several joint configurations manually manipulated by a human; (ii) a current state estimation step, which estimates the object’s current joint angle by optimizing an appropriate loss function; and (iii) a text-guided manipulation step, which determines the target joint angle based on a given text prompt and executes the corresponding robot manipulation. As shown on the left of Figure 8, ScrewSplat accurately reconstructs the shape and kinematic structure of real-world objects – including even a translucent drawer. A well-trained ScrewSplat further enables precise estimation of the current joint angle and facilitates successful text-guided object manipulation, as shown in the middle and right of Figure 8, respectively. Further details for real-world text-guided manipulation are provided in Appendix C.2 and D.3, respectively.
@@ Acknowledgments
S. Kim, J. Ha, Y. H. Kim, and F. C. Park were supported in part by IITP-MSIT under Grant RS-2021-II212068 (SNU AI Innovation Hub), in part by IITP-MSIT under Grant 2022-220480 and Grant RS-2022-II220480 (Training and Inference Methods for Goal Oriented AI Agents), in part by IITP-MSIT under Grant RS-2024-00436680 (Collaborative Research Projects with Microsoft Research) through Global Research Support Program in the Digital Field program, in part by KIAT under Grant P0020536 (HRD Program for Industrial Innovation), in part by SRRC NRF under Grant RS-2023-00208052, in part by SNU-IPAI, in part by SNU-AIIS, in part by SNU-IAMD, in part by SNU Institute for Engineering Research, in part by Hyundai Motor Company and Kia, and in part by Microsoft Research Asia.
@@ A.1 Articulated Object Recognition Using Supervised Learning
There has been considerable interest in recognizing articulated objects within a supervised learning framework, where deep neural networks are trained to predict articulation parameters (e.g., joint axes and joint angles) directly from raw visual input [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. Some of these methods also aim to simultaneously reconstruct the part-aware geometry of objects [5, 8, 10], which aligns with our objectives. While the problem of 3D recognition of rigid objects is well studied and has numerous established solutions [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], the recognition of articulated objects is significantly more challenging, as it requires inferring not only object geometry but also the underlying kinematic structure. Consequently, it remains an open problem.
Early studies often addressed articulated object recognition problem by employing category-driven approaches, using category information to assist in identifying the kinematic structure, and testing on unseen objects belonging to known categories [1, 4, 7, 44]. More recently, to handle a broader range of objects beyond those in known categories, category-agnostic recognition approaches have been proposed. These methods primarily aim to reconstruct part-level geometry and kinematic structures, which closely align with our objective. For example, CARTO [10] predicts the geometry as a defomable signed distance function (SDF) representations – similar to the category-specific A-SDF [4] but removes the reliance on object categories. Ditto [8] predicts articulable neural occupancy fields to generate digital twins of articulated objects from point cloud observations under two different articulation states. Articulate-Anything [45], a recent approach, integrates a vision-language foundation model to recognize part-aware geometries and kinematic structures, and incorporates interactable digital twins into simulators for sim-to-real robot learning. They demonstrate generalizability to unseen objects within similar categories, and even to unseen categories through foundation models. However, they inherently struggle to generalize to objects that differ significantly from the training categories.
@@ A.2 Articulated Object Recognition Using Per-object Optimization Methods
Several works have attempted to recognize articulated objects by directly fitting 3D representations and kinematic structures to observations without any supervision. Since these methods typically perform optimization for each object individually, they are often referred to as per-object optimization methods. A representative early work in per-object optimization methods is PARIS [16], which presents a method based on neural radiance fields (NeRF). Specifically, PARIS defines separate radiance fields for movable and static parts, performing joint rendering and optimization to achieve part-level reconstruction and articulation discovery. However, this approach is only applicable to single-joint objects, i.e., articulated objects with only one movable part, with known joint types.